OpenStudy (anonymous):

Solve V = π r2h for h

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

a = bx => x = a/b

OpenStudy (jamesj):

I suspect you mean \( V = \pi r^2 h \). And now look at what GT just showed you.

OpenStudy (anonymous):

V = π r2h /π r2h /π r2h V = h /π r2h